# algorithmic

### v2

@@ Appendix C Dependent censoring and copulas
Dependent censoring is a type of informative censoring where the dependence between the censoring time and the survival time is not explained by any observable covariate (Emura and Chen, 2018, Ch. 1). In other words, dependent censoring is a consequence of some residual dependence – i.e., when covariates influencing both survival time and censoring time are ignored or not included in the analysis. A copula is a function that links R.V.s by specifying their dependence structure (Emura and Chen, 2018, Ch. 1), thus it can be applied to introduce a dependence between the event and the censoring distribution.
Formally, we write $C(u_{1},\ldots,u_{d}):[0,1]^{d}\rightarrow[0,1]$, which is a $d$-dimensional copula (i.e., where $d$ is the number of variables) with $u_{1},\ldots,u_{d}$ uniform marginal probabilities. For simplicity, we assume the single-event case, where $T_{E}$ is the event time, $T_{C}$ is the censoring time, $\bm{x}$ is a vector of covariates, and $S_{T_{E}}(t\mid\bm{x})=\text{Pr}(T_{E}>t_{e}\mid\bm{x})$ and $S_{T_{C}}(u\mid\bm{x})=\text{Pr}(T_{C}>t_{c}\mid\bm{x})$ are the marginal survival functions given $\bm{x}$. We can then define a bivariate survival copula $C_{\theta}$ that describes the degree of dependence between $T_{E}$ and $T_{C}$ (Emura and Chen, 2018, Ch. 3):
@@
If $C_{\theta}$ satisfies certain mathematical conditions, then Eq. (18) is a valid survival function (Emura and Chen, 2018, Ch. 3). Kendall’s tau ($\tau$) is a common measure of the dependency between $T_{E}$ and $T_{C}$:
@@
where ($T_{1}$, $C_{1}$) and ($T_{2}$, $C_{2}$) are sampled from the model (18). An Archimedean copula is defined as (Emura and Chen, 2018, Ch. 3):
@@
Appendix D Data generation and preprocessing details
Synthetic: This paper utilizes two (linear and nonlinear) synthetic datasets with two competing events (event and censoring), both based on the semiparametric Weibull distribution as discussed in Foomani et al. (2023). The datasets are characterized by the conditional hazard function and conditional survival probability of an event as follows:
@@ Appendix D Data generation and preprocessing details
where $v$ and $\rho$ represent the shape and scale parameters of the Weibull distribution, respectively. We use $v_{E}=4$, $v_{C}=5$, $\rho_{E}=18$, $\rho_{C}=17$. The function $g_{\Psi}(X)$ represents the risk function, which is a linear function for the linear dataset and a single-layer neural network with ReLU activation for the non-linear dataset. The data generation process begins by initializing a copula with a predefined parameter $\theta$. We then sample $N$ sets of CDF observations from the copula, each containing two CDF probabilities (corresponding to each competing event). From these probabilities, we then calculate the corresponding times using the inverse function of (22). This data generation process has been described in Algorithm 2 in Foomani et al. (2023).
@@
SEER: Surveillance, Epidemiology, and End Results (SEER) Program dataset (Ries et al., 2003) is a comprehensive collection of clinical data in the United States. This dataset, which encompasses about 49% of the U.S. population, includes vital information on patient diagnoses, time until an event occurs, and other relevant details sourced from various registries. We use the raw feature “COD to site recode” to extract the event indicator. In particular, we focus on two types of events: breast cancer and heart disease. Then, we apply some special value processing for some features, based on the guideline provided by Wang and Sun (2022). For categorical features, missing values are imputed using the mode, followed by one-hot encoding. The data is then standardized to ensure consistency in our analyses. The SEER cohort is available for download at www.seer.cancer.gov.
@@
MIMIC-IV: We utilized the MIMIC-IV v2.2 dataset (Johnson et al., 2023), which includes data on 299,712 patients and 431,231 admissions. Preprocessing was conducted using the MIMIC-IV Pipeline method (Gupta et al., 2022), through which we extracted 1,672 static features from the dataset. To focus our analysis, we filtered the dataset to include only 26,236 patients aged 60 and 65 years. For feature refinement, we applied the Unicox feature selection method (Qi et al., 2022; Simon et al., 2003). In the training set, we assessed the statistical significance of each feature by calculating the p-values using lifelines’s (Davidson-Pilon, 2019) CoxPHFitter summary. The top 100 features with the smallest p-values were selected. The event of interest in our study is mortality after hospital admission. The final set of features is included in the source code. The MIMIC-IV v2.2 is available for download at https://physionet.org/content/mimiciv/2.2/.
@@
Rotterdam: The Rotterdam dataset (Royston and Altman, 2013) consists of records of 2,982 primary breast cancer patients included in the Rotterdam tumor bank, with 1,546 of these patients having node-positive disease. The dataset contains 10 covariates, including demographic information, tumor characteristics, and treatment details. Survival time is defined as the duration from primary surgery to the occurrence of either disease recurrence or death from any cause, representing two competing risks. During preprocessing, categorical size groups (’$\leq 20$’, ’$20-50$’, ’$>50$’) were converted to their approximate median values (10, 35, 75, respectively). Additionally, the remaining features were normalized to ensure consistency across the dataset. The Rotterdam dataset is available in at https://rdrr.io/cran/survival/man/rotterdam.html.
@@
PRO-ACT: The PRO-ACT dataset (Atassi et al., 2014) contains over 11,600 clinical records (October 4 2024) from different ALS trials, including demographic, lab, family history, and medical data. In each trial, patients adhere to a schedule of visitations until the study concludes, they become too ill to continue or decide to leave the study. Covariates are recorded at each visitation, including total ALSFRS-R scores, age, gender, region of onset (e.g., limb or bulbar), mean forced vital capacity (FVC) scores, time since diagnosis in days, disease progression rate, and whether the patient is taking Riluzole. For our analyses, we only use information about each patient from their baseline visitation. At each visitation in the dataset, that patient’s disease state is evaluated by a clinician using the ALSFRS-R assessment. We consider four events: “Speech”, “Swallowing”, “Handwriting” and “Walking”. We consider an event to have occurred the first time a patient scores 2 or less in that specific assessment during any follow-up visitation following their baseline visitation. These four events can occur in any order and are not mutually exclusive. There are several types of censoring, including death, being unable to perform the assessment or leaving the study prematurely. We use a maximum follow-up time of 500 days to focus on the early stages of the disease. We consider patients that score 2 or less in any of the four assessments at baseline as having already had the event, indicating that the loss of function occurred before trial entry, and exclude patients these from the study. We exclude patients with no recorded ALSFRS-R history. The PRO-ACT dataset is available for download at https://ncri1.partners.org/ProACT.
@@
EBMT: The EBMT dataset (Rossman et al., 2022) consists of 2,279 patient records who underwent blood or marrow transplantation between 1985 and 1998. The dataset is provided by the European Society for Blood and Marrow Transplantation (EBMT). The dataset focuses on survival after transplantation for patients suffering from blood cancer. We considered five events in our analysis: “Recovery” (Rec), “Adverse Event” (AE), a combination of both (AE and Rec), “Relapse”, and “Death”. These events allowed us to evaluate their individual and combined effects on patient prognosis. The dataset includes six features belong to four baseline covariates: donor-recipient match (where a gender mismatch is defined as a female donor and a male recipient), prophylaxis type, year of transplant, and age at transplant. The year of transplantation was encoded as Boolean variables indicating whether the transplant took place during 1990-1994 and 1995-1998, and the age at the transplantation was classified into Boolean variables indicating whether the age was $\leq 20$ or $\geq 40$ years. The EBMT dataset is available in the PyMSM package at https://github.com/hrossman/pymsm.
@@ Appendix E Model hyperparameters
Table 10 reports the selected hyperparameters for the MENSA model. We use Bayesian optimization (Snoek et al., 2012) to tune the hyperparameters over ten iterations in the validation set, adopting the hyperparameters leading to the lowest log-likelihood loss. Table 11 reports the selected hyperparameters for the baseline models. We configure all baseline models with sensible default hyperparameters by consulting the original work.
@@ Appendix F Implementation details and reproducibility
All experiments were implemented and conducted in Python 3.9 using PyTorch 1.13.1, NumPy 1.24.3 and Pandas 1.5.3. Double precision (fp64) was used for tensor computations. All experiments were conducted on a single workstation with an Intel Core i9-10980XE 3.00GHz CPU, 64GB of memory, and an NVIDIA GeForce RTX 3090 GPU. CUDA 11.7 was used to perform model training on the GPU. All datasets used in this study are publicly available. Instructions on how to reproduce the results are available in the source code repository at https://github.com/thecml/mensa
@@ Appendix G Non-proportional hazards in MENSA
This section addresses how the proposed method overcomes the limitations associated with the proportional hazards assumption (Cox, 1972). We begin by reviewing how a traditional Weibull distribution adheres to the proportional hazards property and then show, intuitively, why MENSA, which uses a mixture of Weibull distributions, does not meet this assumption.
It is well-established that the Weibull distribution satisfies the proportional hazards assumption when its shape parameter, $\eta$, is independent of the covariates $\bm{x}^{(i)}$ (Klein and Moeschberger, 2006). The probability density function (PDF), cumulative distribution function (CDF), survival function, and hazard function for a Weibull distribution are given by:
$f(t;\beta,\eta)$ $=\frac{\eta}{\beta}\left(\frac{t}{\beta}\right)^{\eta-1}\exp\left(-\left(\frac{t}{\beta}\right)^{\eta}\right),$
@@
$S(t;\beta,\eta)$ $=1-F(t;\beta,\eta)=\exp\left(-\left(\frac{t}{\beta}\right)^{\eta}\right),$
$h(t;\beta,\eta)$ $=\frac{f(t;\beta,\eta)}{S(t;\beta,\eta)}=\frac{\eta}{\beta}\left(\frac{t}{\beta}\right)^{\eta-1}.$
@@
$\beta(\bm{x}^{(i)})\ =\ \beta_{0}\exp(-\gamma^{\top}\bm{x}^{(i)}),$
@@
$h(t;\bm{x}^{(i)})$ $=\ \frac{\eta}{\beta(\bm{x}^{(i)})}\left(\frac{t}{\beta(\bm{x}^{(i)})}\right)^{\eta-1}$
@@
$h_{0}(t)\ =\ \frac{\eta}{\beta_{0}^{\eta}}\cdot t^{\eta-1}.$
@@
This formulation shows that the Weibull hazard function is the product of a time-dependent baseline hazard, $h_{0}(t)$, and a time-independent relative hazard, $\exp\left(\eta\gamma^{\top}\bm{x}^{(i)}\right)$. Hence, when the shape parameter $\eta$ is constant, and only the scale parameter $\beta$ varies with the covariates, the Weibull distribution satisfies the proportional hazards assumption. In contrast, both the Deep Survival Machines (DSM) (Nagpal et al., 2021) and the proposed MENSA model, which utilize a mixture of Weibull distributions, do not satisfy this assumption. This can be attributed to two main reasons:
@@
    $h\left(t;\,\beta,\eta,\bm{x}^{(i)}\right)\ =\ \frac{f\left(t;\,\beta,\eta,\bm{x}^{(i)}\right)}{S\left(t;\,\beta,\eta,\bm{x}^{(i)}\right)}\ =\ \frac{\sum_{\psi=1}^{\Psi}g_{\psi}(\bm{x}^{(i)})f_{\psi}\left(t;\,\beta_{\psi}(\bm{x}^{(i)}),\eta_{\psi}(\bm{x}^{(i)})\right)}{1-\sum_{\psi=1}^{\Psi}g_{\psi}(\bm{x}^{(i)})F_{\psi}\left(t;\,\beta_{\psi}(\bm{x}^{(i)}),\eta_{\psi}(\bm{x}^{(i)})\right)}.$

### v3

@@ Appendix C Dependent censoring and copulas
Dependent censoring is a type of informative censoring where the dependence between the censoring time and the survival time is not explained by any observable covariate [8, Ch. 1]. In other words, dependent censoring is a consequence of some residual dependence – i.e., when covariates influencing both survival time and censoring time are ignored or not included in the analysis. A copula is a function that links R.V.s by specifying their dependence structure [8, Ch. 1], thus it can be applied to introduce a dependence between the event and the censoring distribution.
Formally, we write $C(u_{1},\ldots,u_{d}):[0,1]^{d}\rightarrow[0,1]$, which is a $d$-dimensional copula (i.e., where $d$ is the number of variables) with $u_{1},\ldots,u_{d}$ uniform marginal probabilities. For simplicity, we assume the single-event case, where $T_{E}$ is the event time, $T_{C}$ is the censoring time, $\bm{x}$ is a vector of covariates, and $S_{T_{E}}(t\mid\bm{x})=\text{Pr}(T_{E}>t_{e}\mid\bm{x})$ and $S_{T_{C}}(u\mid\bm{x})=\text{Pr}(T_{C}>t_{c}\mid\bm{x})$ are the marginal survival functions given $\bm{x}$. We can then define a bivariate survival copula $C_{\theta}$ that describes the degree of dependence between $T_{E}$ and $T_{C}$ [8, Ch. 3]:
@@
If $C_{\theta}$ satisfies certain mathematical conditions, then Eq. (30) is a valid survival function [8, Ch. 3]. Kendall’s tau ($\tau$) is a common measure of the dependency between $T_{E}$ and $T_{C}$:
@@
where ($T_{1}$, $C_{1}$) and ($T_{2}$, $C_{2}$) are sampled from the model (30). An Archimedean copula is defined as [8, Ch. 3]:
@@
Appendix D Data and preprocessing details
Synthetic: This paper utilizes two (linear and nonlinear) synthetic datasets with two competing events (event and censoring), both based on the semiparametric Weibull distribution as discussed in Foomani et al. [12]. The datasets are characterized by the conditional hazard function and conditional survival probability of an event as follows:
@@ Appendix D Data and preprocessing details
where $v$ and $\rho$ represent the shape and scale parameters of the Weibull distribution, respectively. We use $v_{E}=4$, $v_{C}=5$, $\rho_{E}=18$, $\rho_{C}=17$. The function $g_{\Psi}(X)$ represents the risk function, which is a linear function for the linear dataset and a single-layer neural network with ReLU activation for the non-linear dataset. The data generation process begins by initializing a copula with a predefined parameter $\theta$. We then sample $N$ sets of CDF observations from the copula, each containing two CDF probabilities (corresponding to each competing event). From these probabilities, we then calculate the corresponding times using the inverse function of (34). This data generation process has been described in Algorithm 2 in Foomani et al. [12].
@@
SEER: Surveillance, Epidemiology, and End Results (SEER) Program dataset [43] is a comprehensive collection of clinical data in the United States. This dataset, which encompasses about 49% of the U.S. population, includes vital information on patient diagnoses, time until an event occurs, and other relevant details sourced from various registries. We use the raw feature “COD to site recode” to extract the event indicator. In particular, we focus on two types of events: breast cancer and heart disease. Then, we apply some special value processing for some features, based on the guideline provided by Wang and Sun [54]. For categorical features, missing values are imputed using the mode, followed by one-hot encoding. The data is then standardized to ensure consistency in our analyses. The SEER cohort is available for download at www.seer.cancer.gov.
@@
MIMIC-IV: We utilized the MIMIC-IV v2.2 dataset [21], which includes data on 299,712 patients and 431,231 admissions. Preprocessing was conducted using the MIMIC-IV Pipeline method [15], through which we extracted 1,672 static features from the dataset. To focus our analysis, we filtered the dataset to include only 26,236 patients aged 60 and 65 years. For feature refinement, we applied the Unicox feature selection method [39, 47]. In the training set, we assessed the statistical significance of each feature by calculating the p-values using lifelines’s [7] CoxPHFitter summary. The top 100 features with the smallest p-values were selected. The event of interest in our study is mortality after hospital admission. The final set of features is included in the source code. The MIMIC-IV v2.2 is available for download at https://physionet.org/content/mimiciv/2.2/.
@@
Rotterdam: The Rotterdam dataset [45] consists of records of 2,982 primary breast cancer patients included in the Rotterdam tumor bank, with 1,546 of these patients having node-positive disease. The dataset contains 10 covariates, including demographic information, tumor characteristics, and treatment details. Survival time is defined as the duration from primary surgery to the occurrence of either disease recurrence or death from any cause, representing two competing risks. During preprocessing, categorical size groups (’$\leq 20$’, ’$20-50$’, ’$>50$’) were converted to their approximate median values (10, 35, 75, respectively). Additionally, the remaining features were normalized to ensure consistency across the dataset. The Rotterdam dataset is available at https://rdrr.io/cran/survival/man/rotterdam.html.
@@
PRO-ACT: The PRO-ACT dataset [4] contains over 11,600 clinical records (October 4 2024) from different ALS trials, including demographic, lab, family history, and medical data. In each trial, patients adhere to a schedule of visitations until the study concludes, they become too ill to continue or decide to leave the study. Covariates are recorded at each visitation, including total ALSFRS-R scores, age, gender, region of onset (e.g., limb or bulbar), mean forced vital capacity (FVC) scores, time since diagnosis in days, disease progression rate, and whether the patient is taking Riluzole. For our analyses, we only use information about each patient from their baseline visitation. At each visitation in the dataset, that patient’s disease state is evaluated by a clinician using the ALSFRS-R assessment. We consider four events: “Speech”, “Swallowing”, “Handwriting” and “Walking”. We consider an event to have occurred the first time a patient scores 2 or less in that specific assessment during any follow-up visitation following their baseline visitation. These four events can occur in any order and are not mutually exclusive. There are several types of censoring, including death, being unable to perform the assessment or leaving the study prematurely. We use a maximum follow-up time of 500 days to focus on the early stages of the disease. We consider patients that score 2 or less in any of the four assessments at baseline as having already had the event, indicating that the loss of function occurred before trial entry, and exclude patients these from the study. We exclude patients with no recorded ALSFRS-R history. The PRO-ACT dataset is available for download at https://ncri1.partners.org/ProACT.
@@
EBMT: The EBMT dataset [44] consists of 2,279 patient records who underwent blood or marrow transplantation between 1985 and 1998. The dataset is provided by the European Society for Blood and Marrow Transplantation (EBMT). The dataset focuses on survival after transplantation for patients suffering from blood cancer. We considered five events in our analysis: “Recovery” (Rec), “Adverse Event” (AE), a combination of both (AE and Rec), “Relapse”, and “Death”. These events allowed us to evaluate their individual and combined effects on patient prognosis. The dataset includes six features belong to four baseline covariates: donor-recipient match (where a gender mismatch is defined as a female donor and a male recipient), prophylaxis type, year of transplant, and age at transplant. The year of transplantation was encoded as Boolean variables indicating whether the transplant took place during 1990-1994 and 1995-1998, and the age at the transplantation was classified into Boolean variables indicating whether the age was $\leq 20$ or $\geq 40$ years. The EBMT dataset is available in the PyMSM package at https://github.com/hrossman/pymsm.
@@ Appendix E Model hyperparameters
Table 10 reports the selected hyperparameters for the MENSA model. We use Bayesian optimization [48] to tune the hyperparameters over ten iterations in the validation set, adopting the hyperparameters leading to the lowest log-likelihood loss. Table 11 reports the selected hyperparameters for the baseline models. We configure all baseline models with sensible default hyperparameters by consulting the original work.
@@ Appendix F Implementation details and reproducibility
All experiments were implemented and conducted in Python 3.9 using PyTorch 1.13.1, NumPy 1.24.3 and Pandas 1.5.3. Double precision (fp64) was used for tensor computations. All experiments were conducted on a single workstation with an Intel Core i9-10980XE 3.00GHz CPU, 64GB of memory, and an NVIDIA GeForce RTX 3090 GPU. CUDA 11.7 was used to perform model training on the GPU. All datasets used in this study are publicly available. Instructions on how to reproduce the results are available in the source code repository at https://github.com/thecml/mensa.
@@ Appendix G Non-proportional hazards in MENSA
This section addresses how the proposed method overcomes the limitations associated with the proportional hazards assumption [6]. We begin by reviewing how a traditional Weibull distribution adheres to the proportional hazards property and then show, intuitively, why MENSA, which uses a mixture of Weibull distributions, does not meet this assumption.
It is well-established that the Weibull distribution satisfies the proportional hazards assumption when its shape parameter, $\eta$, is independent of the covariates $\bm{x}^{(i)}$ [27]. The probability density function (PDF), cumulative distribution function (CDF), survival function, and hazard function for a Weibull distribution are given by:
$f(t;\beta,\eta)$ $=\frac{\eta}{\beta}\left(\frac{t}{\beta}\right)^{\eta-1}\exp\left(-\left(\frac{t}{\beta}\right)^{\eta}\right),$ (35)
@@
$S(t;\beta,\eta)$ $=1-F(t;\beta,\eta)=\exp\left(-\left(\frac{t}{\beta}\right)^{\eta}\right),$ (37)
$h(t;\beta,\eta)$ $=\frac{f(t;\beta,\eta)}{S(t;\beta,\eta)}=\frac{\eta}{\beta}\left(\frac{t}{\beta}\right)^{\eta-1}.$ (38)
@@
$\beta(\bm{x}^{(i)})\ =\ \beta_{0}\exp(-\gamma^{\top}\bm{x}^{(i)}),$ (39)
@@
$h(t;\bm{x}^{(i)})$ $=\ \frac{\eta}{\beta(\bm{x}^{(i)})}\left(\frac{t}{\beta(\bm{x}^{(i)})}\right)^{\eta-1}$ (40)
@@
$h_{0}(t)\ =\ \frac{\eta}{\beta_{0}^{\eta}}\cdot t^{\eta-1}.$ (41)
@@
This formulation shows that the Weibull hazard function is the product of a time-dependent baseline hazard, $h_{0}(t)$, and a time-independent relative hazard, $\exp\left(\eta\gamma^{\top}\bm{x}^{(i)}\right)$. Hence, when the shape parameter $\eta$ is constant, and only the scale parameter $\beta$ varies with the covariates, the Weibull distribution satisfies the proportional hazards assumption. In contrast, both the Deep Survival Machines (DSM) [36] and the proposed MENSA model, which utilize a mixture of Weibull distributions, do not satisfy this assumption. This can be attributed to two reasons:
@@
    $h\left(t;\,\beta,\eta,\bm{x}^{(i)}\right)\ =\ \frac{f\left(t;\,\beta,\eta,\bm{x}^{(i)}\right)}{S\left(t;\,\beta,\eta,\bm{x}^{(i)}\right)}\ =\ \frac{\sum_{\psi=1}^{\Psi}g_{\psi}(\bm{x}^{(i)})f_{\psi}\left(t;\,\beta_{\psi}(\bm{x}^{(i)}),\eta_{\psi}(\bm{x}^{(i)})\right)}{1-\sum_{\psi=1}^{\Psi}g_{\psi}(\bm{x}^{(i)})F_{\psi}\left(t;\,\beta_{\psi}(\bm{x}^{(i)}),\eta_{\psi}(\bm{x}^{(i)})\right)}.$ (43)